OpenStudy (anonymous):

A line has a slope of -1 and contains the points (1, 2) and (2, y). The value of y is... Can you show me the steps to solve this...

myininaya (myininaya):

\[-1=\frac{y-2}{2-1} => -1=\frac{y-2}{1} =>-1=y-2 => -1+2=y => 1=y\]

myininaya (myininaya):

i just used the formula for the slope

OpenStudy (anonymous):

thanks